ram father is thrice as old as ram was 2 years ago.five years from now his father age will be 6 year more than twice the ram age.what is ram present ???

Answer by chessace(471)

My computer has enough ram for this.
x=ram present age.
f=fathers present age.
f = 3(x-2)
f+5 = 6 + 2(x+5)
3x-6 + 5 = 6 + 2x+10
x = 17